OpenStudy (anonymous):

Find the x and y intercepts of -2x+4y=4

OpenStudy (anonymous):

you can convert it in intercept form y/a + x/b =1 here 0,a are intercept on y axis b,0 on x axis

OpenStudy (anonymous):

To do this, just set x equal to 0 and find y - this gives the y-intercept. Then set y equal to 0 and find x - this gives the x-intercept. -2(0)+4y=4 4y=4 y=1 So the y-intercept is 1 -2x+4(0)=4 -2x=4 x=-2 So the x-intercept is -2

OpenStudy (anonymous):

where on Gods nation does the b come?

OpenStudy (anonymous):

-2x + 4y =4 divide the eq both lhs and rhs by 4 you get x/-2 + y/1 = 1 so co-ordinates are -2,0 and 0,1

OpenStudy (anonymous):

make the opposite variable your looking for zero then sove for the one your are trying to get like -2x+4*0==4 now solve then you will get the answer for x